A proposed cost-saving device has an installed cost of $735,000. The device will be used in a five-year project but is classified as three-year MACRS property for tax purposes. The required initial net working capital investment is $145,000, the marginal tax rate is 23 percent, and the project discount rate is 11 percent. The device has an estimated Year 5 salvage value of $104,000. What level of pretax cost savings do we require for this project to be profitable? MACRS schedule. (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Initial investment, C0 = installed cost of $735,000 + initial net working capital investment is $145,000 = 735,000 + 145,000 =  880,000

Terminal year cash flows:

  • Since the device would be fully depreciated under MACRS schedule, taxable basis will be zero.
  • Post tax salvage value at the end of year 5 = pre tax salvage value x (1 - Tax rate) = 104,000 x (1 - 23%) =  80,080
  • Release of working capital = 145,000
  • Hence terminal year cash flow = 80,080 + 145,000 =  225,080

Discount rate = R = 11%

PV(TCF) = Present value of terminal year cash flow = Terminal year cash flow / (1 + R)5 = 225,080 / (1 + 11%)5 = 133,574

Depreciation tax shield under MACRS 3 year depreciation schedule:

Asset Cost A          735,000
Year N 1 2 3 4
Depreciation rate d 33.33% 44.45% 14.81% 7.41%
Depreciation D = d x A          244,976       326,708       108,854       54,464
Tax rate T 23%
Depreciation tax shield DTS = D x T            56,344          75,143          25,036       12,527
PV Factor @ 11% (1+R)(-N)            0.9009          0.8116          0.7312       0.6587
PV of Depreciation tax shield DTS x PV Factor            50,761          60,988          18,306          8,252
Total PV of Depreciation tax shield; PV (DTS) Sum of all the items in above row          138,306

Let's now assume that pre tax annual cost saving is S. Hence post tax cost savings will be = S x (1 - 23%) = 0.77S

Hence, the present value of post tax cost savings will be = PV of annuity of 0.77S over five years = 0.77S x Sum of PV factors over five years = 0.77S x [1 - (1+R)-N] / R = 0.77S x [1 - (1 + 11%)-5] / 11% =  2.84584S

NPV equation will be:

NPV = -C0 + PV (DTS) + PV(TCF) + PV of post tax annual cost savings over five years = -880,000 + 138,306 + 133,574 + 2.84584S

For this project to be profitable, NPV = 0

or, -880,000 + 138,306 + 133,574 + 2.84584S = 0

Hence, S = (880,000 - 138,306 - 133,574) / 2.84584 =  213,687.21

Level of pretax cost savings do we require for this project to be profitable = S = $ 213,687.21